OpenStudy (anonymous):

Find the product of (-12) and (3

OpenStudy (anonymous):

This one's easy. I know you can do it :P

OpenStudy (anonymous):

Tell me: What's 12 times 3 with a negative attached to it? Though... I'm guessing you meant to add more.

OpenStudy (anonymous):

its hard:( lol

OpenStudy (anonymous):

\[-12\times 3 = -1\times 12 \times 3 = -1(12\times 3)\]

OpenStudy (anonymous):

Just do 12 times 3 and put a negative in the front.